OpenStudy (anonymous):

Can someone help me find the solutions of: g(x) = 4x^2 + 8x + 2 I know it should be two real irrational solutions because of its discriminant, but I do not know the steps to solve it. PLEASE HELP

OpenStudy (anonymous):

just use the shreedharacharya formula for a quadratic ax^2 +bx+c=0 x=\[[-b \pm(\sqrt{b ^{2}-4ac})]/2a\]

OpenStudy (anonymous):

So should I set it up like this \[-8\pm(\sqrt{32})/2(4) ?\]

OpenStudy (cggurumanjunath):

YES

OpenStudy (cggurumanjunath):